OpenStudy (anonymous):

Look at the two triangular flags below. What is the length of the side BC of triangle ABC?

OpenStudy (anonymous):

sorry here you go

OpenStudy (anonymous):

plz help me

jhonyy9 (jhonyy9):

5x-2=x+6 4x=8 x=2 BC=5x-2 BC=5*2-2 BC=10-2 BC=8

jhonyy9 (jhonyy9):

BC=8 cm